# Identified hadron production at hadron colliders in NNLO QCD

### v2

@@ I Introduction
The theory of the strong interactions, Quantum Chromodynamics (QCD), is a pillar of collider physics. Almost all collider measurements rely on QCD, either as a direct input to measurements, or as part of the interpretation of Standard Model (SM) measurements and searches for physics beyond the SM.
Given QCD’s importance, it is natural to ask how one would test and validate QCD itself in the perturbative regime. To this end, ideally, one would like to compare QCD predictions with measurements of observables that are as little sensitive as possible to the electroweak part of the SM. A prime example is the class of observables based on identified light hadrons like $\pi^{\pm},\pi^{0},p$, etc.
Such observables provide direct access to the non-perturbative aspect of QCD via the so-called parton-to-hadron fragmentation functions (FF). A detailed study of the production of single and multiple identified hadrons allows one to study the FF universality and scaling violations, verify QCD factorization theorems and place quantitative constraints on non-perturbative power-suppressed corrections in collider processes.
Observables based on identified hadrons are also sensitive to perturbative corrections, which are computable in perturbation theory. Existing hadron collider studies have been restricted to next-to-leading order (NLO) in QCD Aversa:1988vb ; Jager:2002xm . Very recently soft-gluon resummation for this class of observables was extended to full next-to-next-to leading logarithmic (NNLL) accuracy Hinderer:2018nkb . Hadron-in-jet observables, also considered in this work, are known through NLO and NLL Arleo:2013tya ; Kaufmann:2019ksh ; Liu:2023fsq ; Caletti:2024xaw .
@@
In this work we initiate a research program which aims to conclusively address some of the above questions. Specifically, we use recent advances in pQCD calculations to produce, for the very first time, QCD predictions for single and multiple identified hadrons at hadron colliders with full next-to-next-to leading order (NNLO) precision. We clarify questions like the level of improvement the NNLO QCD correction brings to these observables, if higher order corrections can remove any remaining discrepancy between QCD and data and, finally, assess the possibilities for future extraction of FF with NNLO precision from hadron collider data.
@@ II Calculating the NNLO correction
$D_{i\to h}$ is a set of non-perturbative fragmentation functions which we take from the literature. Specifically, in this work we use the NNFF Bertone:2017tyb and MAPFF AbdulKhalek:2022laj sets for pion final states, and the NNFF Bertone:2018ecm and JAM Moffat:2021dji FF sets for charged hadron observables. The FFs depend on the factorization scale and partonic fraction $z$. All FF sets used in this work are available as LHAPDF library Buckley:2014ana grids, which incorporate DGLAP evolution. Throughout this work we use the parton distribution set NNPDF3.1 NNPDF:2017mvq , which is implicitly included in the partonic coefficient function $d\hat{\sigma}$. The strong coupling constant $\alpha_{S}$ is taken at NNLL as provided by the pdf’s LHAPDF grids.
@@
In our calculations we set the central values of the factorization, renormalization and fragmentation scales to a common value. This value is chosen individually for each observable. The scale choices we make are specified in section III. Scale variation is performed in the 15 point approach.
@@
[…]
×$\times~{}D_{i\to h_{1}}(z_{1})D_{j\to h_{2}}(z_{2})\,,$
with $p_{i}=p_{1}/z_{1}$ and $p_{j}=p_{2}/z_{2}$.
@@ III LHC Phenomenology
In this work we consider three different observables: the transverse momentum $p_{T}(\pi^{0})$ of a neutral pion, the invariant mass $m(\pi^{0},\pi^{0})$ of two neutral pions, and the ratio $z=p_{T}(h)\cos(\Delta R)/p_{T}({\rm jet})$, with $\Delta R=\sqrt{\Delta\phi^{2}+\Delta\eta^{2}}$, of a charged hadron inside a jet. Such a range of observables is sufficient for demonstrating the general behavior of NNLO QCD corrections in hadron collider observables with identified hadrons. All calculations shown in this work, independently of their perturbative order, use pdf with NNLO accuracy. The accuracy of the FF is also kept fixed at their highest available order, NLO or NNLO, depending on the FF set. For all observables involving neutral pions, the $\pi^{0}$ fragmentation functions are derived by averaging the $\pi^{+}$ and $\pi^{-}$ FFs of the corresponding FF set.
@@
Figure 1: Single inclusive $\pi^{0}$ $p_{T}$ spectrum at LO (green), NLO (blue) and NNLO (red) QCD versus ALICE data. Top figure shows ratio to NLO QCD; the colored bands represent scale uncertainty. The inset magnifies the high-$p_{T}$ region. Bottom figure: two NNLO predictions derived with different FF sets. Shown is a breakdown of pdf, FF and scale uncertainty.
@@
Figure 4: As in fig. 3 but at NNLO and for two different FF sets. The bands for each FF set represents its FF uncertainty.
@@
In fig. 3 we focus our attention on the theoretical prediction’s perturbative convergence through NNLO. The predictions shown in this figure utilize the JAM FF set. The main effect of the inclusion of higher order corrections to this observable is the change in slope of the observed distribution. The scale uncertainty at NLO is larger than at LO which is an artifact of this observable. At NNLO the scale variation already slightly decreases relative to NLO which indicates the onset of perturbative convergence. The NNLO correction is significant, for all $p_{T}({\rm jet})$ slices, while remaining consistent with the NLO scale estimate. The overall behavior of this observable is as expected, and it resembles the related distribution for $b$-flavored hadrons studied in great detail in ref. Czakon:2021ohs .
@@
The two FF sets lead to strikingly different predictions. For intermediate $z$ values, $0.20.06<z<$ 0.2, the two are incompatible within their respective FF uncertainties. At large $z$ value the two are compatible, albeit with very different slopes. At low values of $z$ the two are largely consistent although, again, they exhibit very different slopes. As far as data comparison is concerned, neither prediction can satisfactory describe it. Apart form the intermediate $z$ range, the JAM set is consistent with data, while the NNFF set describe data quite well for low and large $z$ but is very far form data in the intermediate $z$ region. The behavior for different $p_{T}({\rm jet})$ cuts is fairly similar. Noting the very small scale uncertainty of this observable at NNLO, we conclude that with improved future FF sets NNLO QCD has the potential to describe charged hadron data at hadron colliders very precisely.

### v3

# Identified Hadron Production at Hadron Colliders in Next-to-Next-to-Leading-Order QCD

@@ I Introduction
The theory of the strong interactions, Quantum Chromodynamics (QCD), is a pillar of collider physics. Almost all high-energy collider measurements rely on QCD, either as a direct input to measurements, or as part of the interpretation of Standard Model (SM) measurements and searches for physics beyond the SM.
Given QCD’s importance, it is natural to ask how one would test and validate QCD itself. To this end, ideally, one would like to compare QCD predictions with measurements of observables that are as little sensitive as possible to the electroweak part of the SM. A prime example is the class of observables based on identified light hadrons like $\pi^{\pm},\pi^{0},p$, etc.
Observables based on identified hadrons are sensitive to perturbative corrections, which are computable in perturbation theory. Such observables also provide direct access to the non-perturbative aspect of QCD via the so-called parton-to-hadron fragmentation functions (FF). A detailed study of the production of single and multiple identified hadrons allows one to study the FF universality and scaling violations, verify QCD factorization theorems and place quantitative constraints on non-perturbative power-suppressed corrections in collider processes.
Existing hadron collider studies have been restricted to next-to-leading order (NLO) in QCD Aversa:1988vb ; Jager:2002xm . Very recently soft-gluon resummation for this class of observables was extended to full next-to-next-to leading logarithmic (NNLL) accuracy Hinderer:2018nkb . Hadron-in-jet observables, also considered in this work, are known through NLO and NLL Arleo:2013tya ; Kaufmann:2019ksh ; Liu:2023fsq ; Caletti:2024xaw .
@@
In this work we initiate a research program which aims to conclusively address some of the above questions. Specifically, we use recent advances in perturbative QCD calculations to produce, for the very first time, QCD predictions for single and multiple identified hadrons at hadron colliders with full next-to-next-to leading order (NNLO) precision. We clarify questions like the level of improvement the NNLO QCD correction brings to these observables, if higher order corrections can remove any remaining discrepancy between QCD and data and, finally, assess the possibilities for future extraction of FF with NNLO precision from hadron collider data.
@@ II Calculating the NNLO correction
$D_{i\to h}$ is a set of non-perturbative fragmentation functions which we take from the literature. Specifically, in this work we use the NNFF Bertone:2017tyb and MAPFF AbdulKhalek:2022laj sets for pion final states, and the NNFF Bertone:2018ecm and JAM Moffat:2021dji FF sets for charged hadron observables. The FFs depend on the factorization scale and partonic fraction $z$. All FF sets used in this work are available as LHAPDF library Buckley:2014ana grids, which incorporate DGLAP evolution. Throughout this work we use the parton distribution function (PDF) set NNPDF3.1 NNPDF:2017mvq , which is implicitly included in the partonic coefficient function $d\hat{\sigma}$. The strong coupling constant $\alpha_{S}$ is taken at NNLL as provided by the PDFs LHAPDF grids.
@@
In our calculations we set the central values of the factorization, renormalization and fragmentation scales to a common value. This value is chosen individually for each observable. The scale choices we make are specified in section III. Scale variation is performed in the 15 point approach, where each scale is varied independently by a factor of two, omitting extreme variations where two of the scales differ by more than a factor of two.
@@
$d\sigma_{pp\to h_{1}h_{2}}(p_{1},p_{2})$ $=$ $\sum_{i,j}\int dz_{1}\int dz_{2}~d\hat{\sigma}_{pp\to ij}\left(\frac{p_{1}}{z_{1}},\frac{p_{2}}{z_{2}}\right)$ (2)
$\times~D_{i\to h_{1}}(z_{1})D_{j\to h_{2}}(z_{2})\,$,
with $p_{i}=p_{1}/z_{1}$ and $p_{j}=p_{2}/z_{2}$. Note that this formula is valid as long as the invariant mass of the two hadrons is sufficiently large to suppress contributions from a single parton fragmenting to two hadrons.
@@ III LHC Phenomenology
In this work we consider three different observables: the transverse momentum $p_{T}(\pi^{0})$ of a neutral pion, the invariant mass $m(\pi^{0},\pi^{0})$ of two neutral pions, and the ratio $z=p_{T}(h)\cos(\Delta R)/p_{T}({\rm jet})$ of a charged hadron inside a jet, where $\Delta R=\sqrt{\Delta\phi^{2}+\Delta\eta^{2}}$ and $\Delta\phi$ and $\Delta\eta$ are the differences in azimuthal angle and pseudorapidity, respectively, between the identified hadron and jet momenta. Such a range of observables is sufficient for demonstrating the general behavior of NNLO QCD corrections in hadron collider observables with identified hadrons. All calculations shown in this work, independently of their perturbative order, use PDFs with NNLO accuracy. The accuracy of the FF is also kept fixed at their highest available order, NLO or NNLO, depending on the FF set. For all observables involving neutral pions, the $\pi^{0}$ fragmentation functions are derived by averaging the $\pi^{+}$ and $\pi^{-}$ FFs of the corresponding FF set. This assumption is often made by FF fitting collaborations and has been found, e.g. in ref. Gao:2025bko , to be valid to very good approximation.
@@
Figure 1: Single inclusive $\pi^{0}$ $p_{T}$ spectrum at LO (green), NLO (blue) and NNLO (red) QCD versus ALICE data. Top figure shows ratio to NLO QCD; the colored bands represent scale uncertainty. The inset magnifies the high-$p_{T}$ region. Bottom figure: two NNLO predictions derived with different FF sets. Shown is a breakdown of PDF, $F$ and scale uncertainty.
@@
Figure 4: As in fig. 3 but at NNLO and for two different FF sets. The bands for each FF set represents its $F$ uncertainty.
@@
In fig. 3 we focus our attention on the theoretical prediction’s perturbative convergence through NNLO. The predictions shown in this figure utilize the JAM FF set. The main effect of the inclusion of higher order corrections to this observable is the change in slope of the observed distribution. The scale uncertainty at NLO is larger than at LO, because the LO scale uncertainty is artificially small. This is typical of normalized distributions, where the scale uncertainty mostly drops out in the ratio at LO. At NNLO the scale variation already slightly decreases relative to NLO which indicates the onset of perturbative convergence. The NNLO correction is significant, for all $p_{T}({\rm jet})$ slices, while remaining consistent with the NLO scale estimate. The overall behavior of this observable is as expected, and it resembles the related distribution for $b$-flavored hadrons studied in great detail in ref. Czakon:2021ohs .
@@
The two FF sets lead to strikingly different predictions. For intermediate $z$ values, $0.06<z<0.2$, the two are incompatible within their respective FF uncertainties. At large $z$ value the two are compatible, albeit with very different slopes. At low values of $z$ the two are largely consistent although, again, they exhibit very different slopes. As far as data comparison is concerned, neither prediction can satisfactory describe it. Apart form the intermediate $z$ range, the JAM set is consistent with data, while the NNFF set describe data quite well for low and large $z$ but is very far form data in the intermediate $z$ region. The behavior for different $p_{T}({\rm jet})$ cuts is fairly similar. Noting the very small scale uncertainty of this observable at NNLO, we conclude that with improved future FF sets NNLO QCD has the potential to describe charged hadron data at hadron colliders very precisely.
All predictions presented above are available online from ref. website .